### v1

# Prethermalization in Fermi-Pasta-Ulam-Tsingou chains

Abstract
The observation of the Fermi-Pasta-Ulam-Tsingou (FPUT) paradox, namely the lack of equipartition in the evolution of a normal mode in a nonlinear chain on unexpectedly long times, is arguably the most famous numerical experiment in the history of physics. Since seventy years after its publication, most studies in FPUT chains still focus on long wavelength initial states similar to the original paper. It is shown here that all characteristic features of the FPUT paradox are rendered even more striking if short(er) wavelength modes are evolved instead. Since not every normal mode leads to equipartition, we also provide a simple technique to predict which modes, and in what order, are excited starting from an initial mode (root) in $\alpha$-FPUT chains. The excitation sequences associated with a root are then shown to spread energy at different speeds, leading to prethermalization regimes that become longer as a function of mode excitation number. This effect is visible in observables such as mode energies and spectral entropies and, surprisingly, also in the time evolution of invariant quantities such as Lyapunov times and Kolmogorov-Sinai entropies. Our findings generalize the original FPUT experiment, provide an original look at the paradox’s source, and enrich the vast literature dedicated to studying equipartition in classical many-body systems.
thermalization slowing-down, prethermalization, Lyapunov spectrum, ergodic hypothesis
I Introduction
The numerical experiment of Fermi, Pasta, Ulam and Tsingou (FPUT) [1] was simultaneously the first computational approach in the study of out-of-equilibrium many-body systems and a source of considerable headache shared by generations of physicists in the past 70 years [1, 2, 3, 4]. A proper understanding of FPUT’s findings, which contradicted standard knowledge at the time and were deemed a paradox, relies heavily on modern mathematical and physical tools unavailable to the authors at that time, e.g. a mature formulation of Kolmogorov-Arnol’d-Moser (KAM) theory [5, 6, 7], resonance overlap criteria [8] and Birkhoff-Gustavson normalization [9, 10, 11, 12, 13], and also on the power of modern computers. Being at the intersection of several fields, it is not surprising that this topic has attracted the attention of so many researchers working on nonlinear dynamical systems and statistical mechanics.
The original FPUT experiment consists in evolving the fundamental mode of a harmonic chain in a weakly perturbed, nonintegrable system of which this mode is no longer a stationary state. More specifically, the authors studied the evolution of the first normal mode of a discretized string under the dynamics generated by the Hamiltonian
$H(\boldsymbol{p},\boldsymbol{x})=\sum_{j=0}^{N}\left[\frac{p_{j}^{2}}{2}+\frac{(x_{j+1}-x_{j})^{2}}{2}+\frac{\alpha(x_{j+1}-x_{j})^{3}}{3}\right]\,,$ (1)
where $p_{j}$ and $x_{j}$ are the momentum and position of each discretized mass in the chain, $\alpha{=}1/4$ and fixed boundary conditions were used, namely $x_{0}{=}x_{N+1}=0$. A very small energy density was chosen for the initial state, i.e. $h{=}H/N\ll 1$, limiting the size of the cubic term and rendering the system a very weak perturbation of the integrable harmonic chain corresponding to setting $\alpha{=}0$. Even for such a weak perturbation the dynamics under (1) ends up being chaotic, justifying the original FPUT expectation that the energy of the initially excited mode would be shared among all other normal modes in a finite (possibly short) time. However, what was verified is that not only energy equipartition was not reached for very long times, but that the energy periodically returned to the initial state and formed recurrences, with a large fraction of modes being barely excited during evolution.
Nowadays, the intersection of a plethora of studies provides a rather converged explanation for the originally observed apparent lack of energy equipartition in finite, weakly anharmonic systems (for reviews, see [2, 14, 15, 16, 4, 17]), although speculations on whether or not the original FPUT setup would eventually reach equilibrium have become a prominent topic in metaphysics. First of all, it is important to note that the energy density chosen by FPUT was extremely small; indeed, a subsequent investigation by Chirikov and Izrailev showed that larger energy densities did result in energy equipartition [18]. Secondly, KAM theory was later proved to be applicable to many types of FPUT-like systems, showing that there is indeed a bound in energy density below which equipartition will never take place [19, 20, 21, 22]. Thirdly, and most relevant to this manuscript, is the fact that the initial state chosen by FPUT is a highly atypical initial state, being a stationary solution of a nearby harmonic chain [23]. Even more, such an atypical initial state is also close to exact stationary time-periodic states coined q-breathers of the anharmonic system under consideration [24, 25]. This strongly contrasts with typical initial conditions, which when expressed in the normal mode basis are combinations of all of its elements. Since such modes form the canonical action variables of the harmonic chain, perturbing a typical initial condition involves an instantaneous coupling of all actions, while for the atypical FPUT initial state only a single action is perturbed. It is therefore only logical that time evolution would be atypical in the original FPUT study, given the atypicality already present in its chosen initial state.
Atypical initial states are, moreover, not all built the same. Their time evolution behaves very differently depending on which initial mode, or root, is excited. The dynamics of roots is characterized by excitation sequences that spread in mode space and, depending on relatively simple algebraic rules, completely or only partially fill it. The case of complete filling is associated to thermal roots, namely states for which time-dependent observables will generally reach equipartition; If the root only excites a subset of modes, known as its bush [26], then equipartition does not take place and the corresponding root is deemed nonthermal. A particularly striking example of nonthermal roots is given by q-breathers, namely families of exact time-periodic orbits which localize in mode space and reduce to the root mode upon tuning the anharmonicity down to zero [24, 25]. Such states show that specific normal modes and their perturbations can also be stationary states of the perturbed chain and provide remarkable counterexamples to the expectation that observables in nonlinear, classical many-body systems necessarily undergo some form of equilibration. They also show that ergodicity-breaking behavior can be observed not only as a function of decreasing energy density, as in the original FPUT experiment, but also as a function of the initial state, as we shall explore here.
This manuscript is devoted to investigating the properties of atypical initial conditions (normal modes) in FPUT chains, and also comparing them to the ones of typical (random) ones. We provide explicit algebraic rules to write mode excitation sequences as a function of the initial root, together with simple criteria to know if such root is thermal or not. We then numerically investigate the time evolution of thermal roots of several different excitation numbers, showing that the paradox associated to the fundamental mode is substantially maximized for higher excited roots, e.g. the prethermal regime and its metastable plateaus become increasingly longer as a function of root excitation number for a fixed energy density. This is shown both by a direct tracking of mode energies as a function of time and by computing equipartition times from spectral entropies.
The impact of root excitation number is not restricted to quantities that depend on the initial state. Indeed, albeit Lyapunov times and Kolmogorov-Sinai entropies being invariant with respect to a large number of transformations and being constant in an ergodic system regardless of the initial state used for computing them [27], we show here that their transient properties are strikingly different from those of typical initial states and clearly reflect the fact that the system is trapped in phase space. More specifically, both the maximal Lyapunov exponent (the inverse of the Lyapunov time) and the Kolmogorov-Sinai entropy show a dip in their time evolution when computed from roots, which is absent if the initial state chosen is typical. Moreover, the duration and depth of this dip increases with excitation number, showing that excited modes remain trapped for even longer in near-integrable portions of phase space. An explanation to this phenomenon is provided by the previously computed excitation sequences of roots, which show that equipartition is fastest for states which spread energy efficiently among modes. The fastest spreader in this sense is always the fundamental root, with the equipartition time slowly increasing with excitation number. Evidently, nonthermal roots are extremely bad spreaders since a portion of mode space is never filled.
The paper is structured as follows. Sec. II discusses the model, which is entirely focused in (1), introducing normal modes, spectral entropies and Lyapunov data which will be thoroughly employed in the rest of the paper. Sec. III presents the rules for computing excitation sequences for roots, which are then evolved in an FPUT chain with $N{=}63$ masses in Sec. IV. A discussion of our findings can be found in Sec. V, together with a conclusion in Sec. VI. An appendix is also included providing examples of pathological numerical behavior due to bush instability.
II Model and observables
We restrict our analysis to cubic FPUT chains, also known as the $\alpha$-FPUT model, with Hamiltonian (1). In the following we recall how to rewrite the FPUT Hamiltonian in normal mode (phonon) basis, and introduce quantities such as mode energies, spectral entropies and equipartition times. We then describe how to obtain important invariant quantities that will be used in the following sections, such as Lyapunov times and Kolmogorov-Sinai entropies.
II.1 Eigenmode basis
Since (1) has $N$ degrees of freedom its harmonic limit has a set of $N$ eigenmodes, denoted $\{P_{j},Q_{j}\}_{j}$, with which we can rewrite the position and momentum in (1) at any time $t$ as
$\begin{pmatrix}p_{j}(t)\\ x_{j}(t)\end{pmatrix}=\sqrt{\frac{2}{N+1}}\sum_{n=1}^{N}\begin{pmatrix}P_{j}(t)\\ Q_{j}(t)\end{pmatrix}\sin\left(\frac{\pi jn}{N+1}\right)\,,$ (2)
each mode having eigenfrequency and eigenenergy
$\omega_{j}{=}2\sin\left[\frac{\pi j}{2(N+1)}\right]\,,\quad E_{j}=\frac{P_{j}^{2}+\omega_{j}^{2}Q_{j}^{2}}{2}\quad.$ (3)
In these canonical coordinates, the Hamiltonian (1) takes the form [19]
$H(\boldsymbol{P},\boldsymbol{Q})=\frac{1}{2}\sum_{j=1}^{N}\big{(}P_{j}^{2}+\omega_{j}^{2}Q_{j}^{2}\big{)}$
$\quad+\frac{\alpha}{3\sqrt{2(N+1)}}\sum_{i,j,k=1}^{N}B_{i,j,k}\,\omega_{i}\,\omega_{j}\,\omega_{k}\,Q_{i}Q_{j}Q_{k}\,,$ (4)
where [28]
$B_{i,j,k}=\delta_{i+j,k}+\delta_{k+i,j}+\delta_{j+k,i}-\delta_{i+j+k,2(N+1)}\,.$ (5)
Note that $B_{i,j,k}$ is invariant with respect to permutations of its indices.
II.2 Spectral entropy
Eq. (1) is much easier to deal with, numerically speaking, than (II.1). However, we are interested in several properties that need to be computed in the eigenmode basis, such as the eigenenergies themselves and the [normalized] spectral entropy
$\eta(t)=\frac{S(t)-\log N}{S(0)-\log N}\,,$ (6)
where the Shannon entropy and normalized eigenenergy are given, respectively, by
$S(t)=-\sum_{j=1}^{N}\rho_{j}(t)\log\rho_{j}(t)\,,\quad\rho_{j}(t)=\frac{E_{j}}{\sum_{j=1}^{N}E_{j}}\,.$ (7)
Spectral entropy has been used several times in order to quantify equipartition in nonlinear systems [23, 4, 29, 30]. The logic for employing it is simple: If ideal equipartition is reached, $\rho_{j}(t)$ will be stationary and the same for every mode $j$, freezing $\eta(t)$ at some value $\langle\eta\rangle$. If one assumes a Gibbs distribution at equilibrium, $\langle\eta\rangle$ can be calculated analytically as a function of the Shannon entropy of the initial state, being given by
$\langle\eta\rangle=\frac{1-\gamma}{\log N-S(0)}\,,$ (8)
where $\gamma{\approx}0.5772$ is the Euler constant [31]. Since $\eta(0){=}1$ for any initial state, one expects $\eta(t)$ to decrease and follow a complicated and possibly highly oscillatory evolution until eventually hitting its stationary value. Once this happens the dynamics cannot, at least on average, depart significantly from thermal oscillations around $\langle\eta\rangle$. It is then natural to consider that equipartition is reached once $\eta(t)$ hits its equilibrium value for the first time [30]. Such first-passage time furnishes our definition of equipartition time, $\tau_{\mathrm{eq}}$, and is appealing because it does not require any ad hoc assumptions.
II.3 Lyapunov invariants and Kolmogorov-Sinai entropy
Equipartition time as defined in Sec. II.2 is, evidently, attached to the observables being monitored, namely the mode energies. Different observables will generally reach equipartition at different times (if at all), such that observable equipartition does not actually measure properties intrinsic to the system at hand. Nevertheless, since mode energies are conserved in the unperturbed harmonic chain, one knows well how these observables should behave upon approaching an integrable limit, i.e., equipartition should slow down and eventually stop for sufficiently small energy densities.
There is, however, an observable-independent time scale that is completely intrinsic to the system and should not depend on the initial state, given by the Lyapunov time (the inverse of the maximal Lyapunov exponent). Since, like the maximal Lyapunov exponent (MLE), all exponents in the Lyapunov spectrum (LS) are invariant with respect to homeomorphisms [27], the LS provides a valuable set of time scales related to the average “pull” along each 1-dimensional submanifold that forms the hyperbolic skeleton of the system’s dynamics in phase space. The information in the LS can also be condensed in the form of a scalar, the Kolmogorov-Sinai entropy, which by Pesin’s theorem can be obtained as the sum of all positive exponents in the LS [32].
In order to compute the LS we employ the well-known prescription of [33, 34], which amounts to computing the spectrum of the symmetric matrix
$\Gamma(\boldsymbol{p}_{0},\boldsymbol{x}_{0})=\lim_{t\to\infty}\left[\mathcal{M}^{T}(\boldsymbol{p}_{0},\boldsymbol{x}_{0};t)\mathcal{M}(\boldsymbol{p}_{0},\boldsymbol{x}_{0};t)\right]^{1/2t}\,.$ (9)
In the above $\mathcal{M}$ represents the system’s stability matrix, which is obtained by solving Hamilton’s equations in tangent space:
$\frac{\mathrm{d}\mathcal{M}(\boldsymbol{p},\boldsymbol{x};t)}{\mathrm{d}t}=\mathcal{J}\mathrm{Hess}\left[H(\boldsymbol{p},\boldsymbol{x})\right]\mathcal{M}(\boldsymbol{p},\boldsymbol{x};t)\,,$ (10)
where $\mathrm{Hess}$ is the Hessian with respect to $(\boldsymbol{p},\boldsymbol{x})$. Since the stability matrix has to be computed along a trajectory, the above equation is solved in parallel with Hamilton’s equations for an initial phase-space point $(\boldsymbol{p}_{0},\boldsymbol{x}_{0})$ until a final time that is long enough to result in a converged LS. During this evolution, one must also perform QR-diagonalizations of $\mathcal{M}$ in order to avoid numerical denegenacies, as is well-known and described in, e.g., [34].
As can be seen in (9), the stability matrix will generally depend on the trajectory along which it is calculated, i.e., on the initial phase-space point $(\boldsymbol{p}_{0},\boldsymbol{x}_{0})$. However, for a sufficiently large number of degrees of freedom and energy density, phase space will consist of a single metrically intransitive set that is equally accessible to all trajectories in the system [35, 36]. Equivalently, one can formulate this previous property as stating that, for sufficiently large $N$ and $h{=}E/N$, the system lies outside the KAM regime in which phase space can be split into a chaotic web and pockets of near-integrability. Thus, since all trajectories are “free to roam” around the entire accessible phase space, the LS is expected to be independent of the trajectory along which it is computed. This independence is a stronger evidence that the system at hand is ergodic than computing observables, and can be ascertained directly from the verification that the KS entropy is independent of the trajectory along which it is computed.
III Excitation sequences
We start this section by writing a very abridged version of Newton’s second law for a given mode $i$, obtained from (II.1):
$\ddot{Q}_{i}+\omega_{i}^{2}Q_{i}\sim\sum_{j,k=1}^{N}B_{i,j,k}\,x_{j}Q_{k}\,.$ (11)
Thus, mode $Q_{i}$ is excited either by a lone mode $x_{j}$ or a mixture of modes $x_{j}$ and $Q_{k}$. Excitations from single modes have $j=k$ and the coupling tensor assumes the form
$B_{i,j,j}$ $=\delta_{i+j,j}+\delta_{j+i,j}+\delta_{2j,i}-\delta_{i+2j,2(N+1)}$
$\Longrightarrow\quad B_{i,j,j}$ $=\delta_{i,2j}-\delta_{i,2(N+1)-2j}\,,$ (12)
where the first two terms vanished because $j$ and $i$ are never zero. Now, for mixed excitations we have $j\neq k$, so
$B_{i,j,k}$ $=\delta_{i+j,k}+\delta_{k+i,j}+\delta_{j+k,i}-\delta_{i+j+k,2(N+1)}$
$\Longrightarrow\quad B_{i,j,k}$ $=\delta_{i,k+j}+\delta_{i,|k-j|}-\delta_{i,2(N+1)-(j+k)}\,.$ (13)
The above expression is composed of: An ascending term, namely $\delta_{i,k+j}$, which dictates which higher-order mode to excite starting from $j$ and $k$; A right reflection term, $\delta_{i,2(N+1)-(j+k)}$, which dictates which term to excite if $j+k>(N+1)$, i.e. if the next mode in the ascending sequence is too large and leaves mode space; And a descending term, $\delta_{i,|k-j|}$, which propagates the excitation backward in mode space and is also responsible for controlling left reflections, which take place when the next excited mode in a descending sequence is smaller than zero.
III.1 First excitation cycle and q-breathers
Initializing the system in a root $Q_{r}$ means setting $x_{j}{=}0$ unless $j{=}r$. The first excitation cycle, therefore, starts with a lone mode $Q_{r}$, which by (III) excites modes $i=2r$ and $i=2(N-r+1)$. If the first condition is met, then the second delta will be one iff $r=(N+1)/2$, with $N+1$ even. Thus, if the chain has an odd number of sites and the root chosen is $Q_{(N+1)/2}$, then all coefficients in the coupling tensor vanish and the energy remains forever localized in the root, i.e. this mode is a q-breather.
Now, assuming that the second Kronecker delta in (III) vanishes, i.e. that $i\neq 2(N-r+1)$, then the only mode that is excited starting from $Q_{r}$ is $Q_{2r}$. This constitutes the first excitation cycle, namely the first run over $j$ and $k$ indices in (11). Note that $Q_{r}$ excites $Q_{2r}$ only if $2(r-1)<N$, since $i$ cannot be negative or zero. This justifies splitting excitation sequences into two types, namely low- and high-mode excitations, depending on whether the root $r$ fulfills $2(r-1)<N$ or not. The only difference is that for low-mode excitations the initial sequence will be ascending, while it is descending in the high-mode case. Although these scenarios are essentially identical, we will assume low-mode excitations here in order for sequences to start in the ascending phase.
III.2 First ascending sequence
The second cycle for low-mode excitations starts with two excited modes, $Q_{r}$ and $Q_{2r}$. Evidently, $Q_{r}$ will re-excite $Q_{2r}$, which on the other hand will hit $Q_{4r}$ due to (III). According to (III) mode $Q_{3r=(2r+r)}$ will also be excited. Thus, the second excitation cycle results in the sequence $\{(Q_{r}),(Q_{2r}),(Q_{3r},Q_{4r})\}$. For the third cycle, $Q_{6r}$ and $Q_{8r}$ will be present due to (III), together with all terms that can be formed from $j+k$ sums as dictated by (III). In this ascending phase, terms created from $|j-k|$ are not important, since any $j$ and $k$ will be multiples of $r$ that were already excited in previous cycles. Thus, by induction, the first ascending sequence starting from a root $Q_{r}$, with $2(r-1)<N$, is simply
$\{(Q_{r}),(Q_{2r}),(Q_{3r},Q_{4r}),(Q_{5r},\dots,Q_{8r})$
$\qquad(Q_{9r},\dots,Q_{16r}),(Q_{17r},\dots,Q_{32r}),\dots\}\,$ (14)
until a reflection takes place in (III), i.e. until $j+k=2(N+1)$ for some $x_{j}$ and $Q_{k}$. The reflection, however, will only excite new modes if $2(N+1)-(j+k)$ is not a multiple of $r$, since otherwise the terms would have been already excited during the ascending sequence. This will be explored in the following.
III.3 First descending sequence
At the end of the first ascending sequence every term is a multiple of the root, $r$. Thus, at the reflection, substituting $j+k=a\,r$ in the last delta of (III) singles out $r=2(N+1)/a$, allowing us to state two important facts:
  1. 1.
    If $r$ divides $2(N+1)$ the first ascending sequence will be reflected into itself (for low modes) or be equivalent to a low mode excitation sequence that reflects into itself (for high modes). This shows, in particular, that $Q_{2}$ is nonthermal for all $N$, since $2(N+1)$ is a always divisible by 2. It also shows that $Q_{1}$ is a trivial thermal root, since it excites all modes in the first ascending sequence for any $N$;
  2. 2.
    The reflection takes place when a multiple of $r$ is larger than $N$ for the first time, i.e. when $x_{j}=Q_{(a+1)r}$. Evidently this mode lies outside mode space and the excited mode is actually $Q_{2(N+1)-(a+1)\,r}$ after the reflection. A consequence is that even roots can never excite odd modes, since for even modes $2(N+1)-(a+1)\,r$ is always even. This shows that $Q_{2r}$ are nonthermal roots for all $r$ and all $N$.
Once a reflection occurs, the selection rule for $|j-k|$ generates a descending sequence starting from $Q_{2(N+1)-(a+1)\,r}$, namely,
$(Q_{2(N+1)-(a+2)\,r},Q_{2(N+1)-(a+3)\,r},\dots)\,.$ (15)
The descent continues in multiples of $r$ until the next mode would exit the chain, i.e. until $02(N+1)-(a+b)\,r<$ 0.
III.4 Full excitation sequences
Once the first descending sequence is over, another reflection takes place, now going from left to right, mediated by $\delta_{i,|j-k|}$ in (III). Substituting the last mode number from the descending sequence in this delta, we see that the excited mode after this reflection is $Q_{2(N+1)-(a+b)\,r+r}$. The second ascending sequence will start from this mode number, once again ascending in steps of $r$ until $2(N+1)-(a+b-c)r>N$. Then, we have another reflection from left to right that excites mode $2(N+1)-[2(N+1)-(a+b-c)r-r]=(a+b-c+1)r$, etc. Instead of writing full abstract sequences, we now provide a few examples showing how easy such sequences are to write in practice.
Figure 1: Evolution of mode energies for the root $Q_{3}$ in an $\alpha$-FPUT chain with (a) $N{=}11$ and (b) $N{=}12$ masses. This root is only thermal in the latter case, while in the former its bush is composed of all multiples of $3$ that are smaller than $11$. The mode excitation sequence in (b) is the same as the one predicted using the coupling coefficients, (III.4).
Let us start with a chain with $N=11$ masses. Since $242(N+1)=$ 24, the only three nontrivial thermal roots in this chain are $Q_{5}$, $Q_{7}$ and $Q_{11}$. The root $Q_{3}$ is evidently nonthermal since it divides 24, but let’s take a look at its excitation sequence. The first ascending sequence is $\{(Q_{3}),(Q_{6}),(Q_{9})\}$, and we stop here because the next term falls outside mode space. The reflection term comes from $i=2(N+1)-(6+9)=9$, which is already excited. This will be now propagated backwards to $Q_{6}$ and $Q_{3}$ accoding to (III), retracing the ascending sequence and never exciting any new mode. Now, for the leftmost reflection, the excited mode is $Q_{|3-3|}$, which does not exist. Thus, the bush of $Q_{3}$ is given by $\{(Q_{3}),(Q_{6}),(Q_{9})\}$. This is verified numerically in Fig. 1(a).
Figure 2: Evolution of mode energies for the first five odd modes, namely: (a) $Q_{1}$; (b) $Q_{3}$; (c) $Q_{5}$; (d) $Q_{7}$; (e) $Q_{9}$; (d) and a random initial condition, all with energy density $h{\approx}0.085$ in an $\alpha$-FPUT chain with $N{=}63$. Line colors correspond to mode frequencies, going from lower (blue) to higher (green), with their equilibrium value $h{=}E/N$ shown as a black dashed line. Equipartition times (orange triangles) are defined as the instant the time-evolving spectral entropy, shown in black in the inset, first attains its equilibrium value $\langle\eta\rangle$, given by the horizontal red line.
Let us now consider what happens to $Q_{3}$ for $N=12$ instead. The first ascending sequence is now updated to $\{(Q_{3}),(Q_{6}),(Q_{9},Q_{12})\}$. The reflection gives $i=2(N+1)-(3+12)=11$, which was not previously excited and is not a multiple of $r=3$. This will then propagate backwards to $11-3=8$, $11-6=5$, etc. The full excitation sequence is then given by
$\nearrow$ $\quad\{(Q_{3}),(Q_{6}),(Q_{9},Q_{12},Q_{11=26-(12+3)}),$
[…]
$\nearrow$ $\quad(Q_{4},Q_{7},Q_{10})\}\,,$ (16)
[…]
As another example, for $Q_{3}$ and $N=13$ we have
$\nearrow$ $\quad\{(Q_{3}),(Q_{6}),(Q_{9},Q_{12},Q_{13=28-(12+3)}),$
[…]
$\nearrow$ $\quad(Q_{2=|1-3|},Q_{5},Q_{8},Q_{11})\}\,.$ (17)
In both (III.4) and (III.4) the root $Q_{3}$ resulted in three excitation sequences: two ascending and one descending. If $Q_{5}$ is chosen as the root, then the full excitation sequence for $N=13$ is
$\nearrow$ $\quad\{(Q_{5}),(Q_{10},Q_{13=28-(10+5)}),$
[…]
$\nearrow$ $\quad(Q_{7},Q_{12},Q_{11=28-(12+5)}),$
[…]
$\nearrow$ $\quad(Q_{4=|1-5|},Q_{9})\}\,.$ (18)
In general, it is easy to see that since every step in the sequence is performed in multiples of $r$, one needs $r$ ascents and descents (or $r-1$ reflections) to cover mode space for a thermal root $Q_{r}$, no matter what $N$ is chosen. Since large $r$ requires more cycles to cover mode space, it means less modes are excited at the end of each descending or ascending sequence. One can once again use the case of $N=13$ as an example: Here, at the end of the first ascending sequence, the root $Q_{1}$ has already excited all modes in the chain, which is a unique property of the fundamental root; $Q_{3}$, however, has only excited 5 of the 13 available modes, and $Q_{5}$ only 3 out of 13. In Sec. IV we will show that this effect strongly impacts any quantity computed from evolving roots.
IV Numerical experiments
Figure 3: Mode energies, $E_{j}$, as a function of mode number, $j$, for roots $Q_{1}$, $Q_{3}$ and $Q_{9}$ for fixed times, with their equilibrium (average) value, $h$, shown as a black dashed line. Each panel is a vertical slice of a panel in Fig. 2 at (a) $t{=}\tau_{\mathrm{eq}}/100$; (b) $t{=}\tau_{\mathrm{eq}}$ and (c) $t{=}10\tau_{\mathrm{eq}}$ (note that $\tau_{\mathrm{eq}}$ is different for each root). The short time of (a) allows for a clear visualization of the energy cascade flowing between modes that are multiples of 1, 3 and 9, as described in Sec. III. At equipartition most of the energies per site have the same order of magnitude, as seen in (b), with stronger mixing still taking place at high energy modes while the lower half of the excitation sequence remains slightly isolated from the other modes. Even after ten equipartition times the nearly homogeneous energy distribution hints at which root was initially excited, as can be seen in the above-average energies of modes $Q_{9}$ and $Q_{18}$ for the root $Q_{9}$ in (c).
In this section we perform computations in the $\alpha$-FPUT chain (1) with $N{=}63$ sites. This value of $N$ is chosen because $1282(N+1)=$ 128 is not divisible by any odd integer other than 1, such that all odd roots are thermal for this chain. Although initial states, final propagation times and energy densities will vary, evolution is always computed using an optimized second order symplectic integrator [37] with step size $\Delta t=0.2$. This results in a maximum relative energy error of around $10^{-3}$ for the largest energy densities chosen, and much smaller for most of them. We will focus on the evolution of random initial conditions and the first five odd roots $Q_{1}$, $Q_{3}$, $Q_{5}$, $Q_{7}$ and $Q_{9}$.
IV.1 Metastable plateaus
Sec. III described how energy can remain localized within a small subset of modes before spreading throughout mode space: The higher excited the root, the smaller the initial subset of excited modes, and also the slower the energy transfer. During this transfer time the dynamics is mostly localized within the excited modes, which essentially “wait” while slowly transferring energy to the ones in the next cycles, forming the metastable plateaus well known in FPUT literature [17, 38]. The longer it takes for the energy to reach the end of the sequence, the longer a metastable plateau survives.
The well-known fact that the metastable plateaus associated to roots become more visible at lower energy densities is unsurprising, since in the limit of zero energy density such modes are normal modes of the associated harmonic lattice. However, the behavior of plateaus for excited roots was not previously addressed. To this end, Fig. 2 shows a comparison of evolving mode energies and spectral entropies for the prototypical $\alpha$-FPUT chain (1) computed starting from the first 5 odd roots in a chain with $N=63$ masses, together with a random initial state. Fig. 2(a) displays the standard mode energy dynamics found in literature: An excitation of $Q_{1}$ monotonically spreads to all other modes and eventually results in energy equipartition, with all $E_{k}$’s converging to the mean energy $E/N=h$ (black dashed line). Metastable plateaus are clearly visible in the prethermal regime, in which a measurement would indicate that the system is not ergodic and cannot be described by equilibrium statistical mechanics. However, such a statement is evidently incorrect, as waiting a bit longer would result in a system that does display statistical behavior.
The approach from an out-of-equilibrium initial state towards a thermal one is most clearly seen in tracking the spectral entropy as a function of time, shown in the insets: The moment it touches the red line, which represents its equilibrium value (8), is where we consider equipartition to have been achieved. The associated equipartition times are displayed in the main plots as an orange triangle and slowly increase as we move from $Q_{1}$ up to $Q_{9}$, with the shortest time corresponding to the random initial condition. Evidently, this latter case is less interesting since the equipartition time depends strongly on the initial sampling, but for roots the structures formed by mode energies as they evolve in time show that, indeed, higher-excited roots isolate an increasingly smaller subset of modes in mode space. Fig. 2(a) also shows a dense subset of modes homogeneously concentrated around the mean, while Fig. 2(e) clearly shows that two modes, namely $Q_{9}$ and $Q_{18}$, remain relatively isolated from the others for much longer times. To see this more clearly, in Fig. 3 we take vertical slices of Fig. 2 at three different fixed times for $Q_{1}$, $Q_{3}$ and $Q_{9}$. At $t{=}\tau_{\mathrm{eq}}$ low frequency modes that are multiples of 3 and 9 can still be seen to have above-average energies for roots $Q_{3}$ and $Q_{9}$. At $t{=}10\tau_{\mathrm{eq}}$ this effect has already dissipated for $Q_{3}$, but the energies of modes $Q_{9}$ and $Q_{18}$ remain slightly above the mean energy when evolving $Q_{9}$, as visible in Fig. 3(c).
Figure 4: Standard deviation of mode energies in Fig. 2 as a function of time for $h{\approx}0.085$. Despite keeping the energy density fixed, the prethermal regime becomes longer as a function of root number.
Evidently, the longer it takes for a subset of modes to reach equipartition, the longer the prethermal regime. This is clearly seen in Fig. 2: The relatively obfuscated metastability associated to the $Q_{1}$-root seen in panel (a) is already maximized in (b), where $Q_{3}$ is evolved, and becomes more and more blatant as higher-excited modes are used as roots. Evidently, the random initial condition in Fig. 2(e) does not display metastability, since here all modes are excited at once and the slow, selection-rule mediated transferring of energy from excited modes to previously unexcited ones does not take place. Since the mean energy is a constant of motion, the prethermal regime’s duration can also be visualized by plotting the standard deviation of $E_{k}(t)$, $\sigma[E_{k}(t)]$, as a function of time, which will be approximately constant while the energy is still strongly localized in a subset of normal modes. This is indeed seen to be true in Fig. 4, where it is also clear that the metastability’s lifetime increases monotonically as a function of root excitation for this energy density.
IV.2 Equipartition and Lyapunov times
Figure 5: Equipartition times computed by evolving the first five odd roots (rainbow colors) and 280 random initial states (tones of magenta and orange) at several different energy densities. These data were obtained from spectral entropies computed up to $t{=}10^{8}$, such that the top left equipartition times, which are of $\mathcal{O}(10^{8})$, are likely slightly overestimated by being of the order of the computation time. The equipartition time for $Q_{1}$ (red diamonds) clearly forms an upper bound that the ones computed from random initial states essentially never cross. The equipartition times computed from excited modes consistently do so, albeit slowly as a function of increasing root number.
Spectral entropy (6) is a function of the initial state, so it is no surprise that the equipartition times computed from it in Fig. 2 depend on it. These times have no direct relation to the duration of prethermalization, except for the fact that they are necessarily longer than the length of metastable plateaus. Since random initial conditions do not, at least generally, undergo prethermal regimes, it is expected that the equipartition times of such typical initial conditions will be shorter than the ones of roots for any energy density chosen. This expectation is confirmed in Fig. 5, where we compare equipartition times for 280 typical conditions with the ones obtained from odd roots as a function of the energy density. Clearly, the former are bounded from above by the latter, which slowly increase as a function of excitation number.
If equipartition times reach the order of the propagation time used to compute them, namely $\mathcal{O}(10^{8})$, convergence can no longer be achieved for these parameters, which is clear in Fig. 5 by the presence of missing data for some values of energy density. However, the “disappearance” of points does not increase fully monotonically with mode excitation number, with some data displaying what could be described as “erratic behavior”. Since dependence on initial conditions is tantamount to ergodicity breaking, investigating quantities that are proved to be invariant in any ergodic system will shed light into the source of the instabilities seen in Fig. 5. The most important of such quantities is the MLE, $\lambda_{1}$, which indicates the presence of chaotic behavior and is not only invariant with respect to diffeomorphisms but also constant for ergodic dynamical systems [27]. From this exponent one can extract the Lyapunov time, $\tau_{1}=1/\lambda_{1}$, which measures the time a vector tangent to a trajectory takes to align with the maximally chaotic, one-dimensional submanifold in the tangent bundle [39, 40]. Since the Lyapunov time is computed from the divergence of two initially close trajectories, it takes a finite time to converge and it is often useful to observe its evolution as a function of time, as done in Fig. 6.
Figure 6: (a) Time-dependent Lyapunov times computed by evolving the first five odd modes at a fixed energy density of $h{\approx}0.085$ up to $t_{f}{=}10^{9}$. Clearly, at this energy density all modes behave similarly and indistinguishably from the random initial states shown in panel (b), for which the propagation time was set to $t_{f}{=}10^{8}$. (c) By lowering the energy density down to $h{\approx}0.033$, roots start to behave differently as a function of time when compared to the random states used in (d). The differences appear to grow with excitation number, resulting in high-order modes such $Q_{9}$ not converging at all within this propagation interval (which is more than enough for all other initial states).
Panels (a) and (b) of Fig. 6 display the evolution of $\tau_{1}(t)$ extracted from propagating odd roots and random initial conditions for an energy density of $h{\approx}0.085$. Not only is convergence achieved for all initial states employed, but the time-dependent portrait of both typical and atypical initial conditions is identical. Upon decreasing the energy density, however, Fig. 6(c) shows that atypical initial conditions start behaving very differently not only with respect to typical ones, but also among themselves. While the Lyapunov times obtained from $Q_{1}$ and $Q_{3}$ are similar to the ones of the random initial conditions in Fig. 6(d), the data obtained from evolving $Q_{5}$ is seen to start forming a “belly” before convergence. Thus, it shows that $Q_{5}$ lies in a region of phase space that takes longer to align with the maximally chaotic direction, similarly to the trapping times described in [23]. The difference is that here we observe longer trapping times for a fixed energy density by changing only the initial conditions.
Figure 7: (a) Lyapunov times computed by evolving the first five odd roots (rainbow colors) at several different energy densities. Propagation time used was $t_{f}{=}10^{9}$, for which $\tau_{1}$’s computed by evolving $Q_{1}$ converge well for all energy densities used. Other modes, however, did not converge for the whole range of energy densities employed here, requiring longer and unpractical propagation times (see text for explanation). (b) Lyapunov times computed by evolving 10 random initial states at the same energy densities as panel (a). Propagation times in this case were only $t_{f}{=}10^{7}$, which is enough to obtain converged results starting from typical initial conditions. Error bars in all plots are the standard deviation of the corresponding $\tau_{1}(t)$ between $t_{f}$ and $t_{f}/10$.
By evolving even higher excited roots such as $Q_{7}$, Fig. 6(c) shows that the “belly” increases when compared to the one of $Q_{5}$, such that $Q_{7}$ is trapped for a time one order of magnitude longer. Extrapolating to $Q_{9}$ this transient trapping regime, which is a manifestation of prethermalization in an invariant quantity and therefore must disappear for sufficiently long times, ends up being longer than the propagation time and does not converge in our computations. The reason for the longer transient times as we go up from $Q_{1}$ to $Q_{9}$ is that the higher the excitation number, the smaller the subspace of mode space in which dynamics is approximately restricted to during the prethermal regime, as discussed in Sec. III. Indeed, Fig. 3 shows that a large fraction of the energy is still concentrated in multiples of the initial excitation number at equipartition time, which as seen by comparing Figs. 6 and 7 is always longer than Lyapunov time when dealing with atypical initial conditions. Thus, as long as the root is a low lying mode (that is, $Q_{r}$ with $2(r-1)<N$), the trapping time will increase with the excitation number. The breather at $Q_{32}$ marks the boundary between low-mode and high-mode excitations, with $Q_{31}$ being the slowest mode to thermalize for $N{=}63$–so slow that there is no point in attempting to obtain data from it, since this is already computationally hard for $Q_{9}$. It is interesting to note that components of the trajectory that “scan” phase space in order to obtain convergence in $\tau_{1}$ are the low-energy ripples that are not multiples of $r$ for a root $Q_{r}$. Evidently, these ripples might not have enough energy to go very far, requiring longer transient times. The energy is spread among all modes most efficiently when evolving $Q_{1}$, which explains why this state behaves essentially as a random state despite its extremely long equipartition time when compared to the latter.
Figure 8: Time-dependent Kolmogorov-Sinai entropy in an $\alpha$-FPUT chain with $N{=}63$ masses computed from roots (rainbow colors) and from a random state (black). Each panel is computed for initial states with decreasing energy densities, $h$, shown in the left corner.
Fig. 7, which shows the final (converged) Lyapunov times as a function of energy density as obtained from typical and atypical initial conditions, is clear evidence that the system we are dealing with is numerically ergodic: The times are the same, no matter what initial conditions are chosen (although the larger error bars and spread for $h{\approx}10^{-2}$ show that $t_{f}{=}10^{9}$ is barely long enough to reach numerical ergodicity for the smallest values of $h$ used here). This does not mean, as discussed before, that all data for roots converges, with the main reason for failure being the presence of extremely long prethermal regimes as seen in Fig. 6.
IV.3 Kolmogorov-Sinai entropies
Given the imprint prethermalization has in a dynamical invariant such as the Lyapunov time, in Fig. 8 we display the time evolution of one of the most meaningful, yet computationally expensive, invariant quantities in dynamical systems, namely the Kolmogorov-Sinai entropy, $\kappa_{\mathrm{KS}}$. This quantity, just like the Lyapunov time and the associated maximal exponent, displays a clear dip when computed by propagating roots, while for a typical initial condition it approaches its converged value from above. Since $\kappa_{\mathrm{KS}}$ is the sum of all positive Lyapunov exponents, what Fig. 8 shows is that the prethermal regime is present in all phase-space directions, not only the maximal one associated to the MLE. Upon decreasing the energy density we also see a tendency of a larger dip for higher-excited modes, in accordance with our expectation that the prethermal regime is more pronounced for higher excitation numbers. We also note that convergence of Kolmogorov-Sinai entropies to the same final value is likely the strongeast display of numerical ergodicity achievable in simulations.
V Discussion
The root chosen in the original FPUT experiment and most subsequent investigations of the FPUT paradox was the fundamental mode. In essence, and as previously shown by comparisons with the nearby and integrable Toda chain [23], the proximity of the fundamental mode to a stationary state of the FPUT chain ends up trapping it in a near-regular portion of phase space for a finite time. We have shown here that higher-excited roots are trapped for even longer than the fundamental mode in such near-regular regions, even though all other parameters in the system are held constant.
Traditionally, the Lyapunov time is considered to characterize alignment along the maximally chaotic direction in a decomposition of the tangent space known as Oseledec splitting, which is covariant with respect to the Hamiltonian flow [40]. Evidently, such decomposition is meaningless in the case of a near-regular trajectory, and it is expected that the transient regime of $\lambda_{1}(t)$ will be different depending on how close the initial state is to a near-integrable region. However, the Lyapunov time itself is an invariant quantity in ergodic systems, such that at the end all initial states must provide the same value for the converged MLE (that is, the asymptotic limit of $\lambda_{1}(t))$. This was verified in Figs. 6 and 7. Thus, the time needed to escape a region of near-integrability is more related to the equipartition time than to Lyapunov time, since we expect mode space to be fully covered only once a trajectory starting in a root becomes sufficiently chaotic. The Lyapunov time itself then carries absolutely no information on prethermalization, although its time-evolution certainly does. This leads to a reinterpretation of $\tau_{1}$ as the time a typical initial condition takes to align with the maximally chaotic covariant direction in tangent space, since atypical ones might not align with it at all. This results in the Lyapunov time being always shorter than the equipartition time of a root, as seen by plotting them together in Fig. 9.
Figure 9: Comparison of Lyapunov times in Fig. 7, where here both panels are superposed, and its corresponding linear fit. The equipartition times for root modes in Fig. 5 are also displayed, together with a linear fit for those of $Q_{1}$, showing that they are orders of magnitude longer than the Lyapunov times for all energy densities.
The only time scale longer than the equipartition time of roots appears to be the convergence time of their time-dependent Kolmogorov-Sinai entropies, $\kappa_{\mathrm{KS}}(t)$. Like $\lambda_{1}(t)$, this quantity will only converge once the trajectory leaves a near-regular region, but now alignment with respect to a large fraction of covariant Lyapunov directions is required instead of only the maximal one. Since several directions are associated to small Lyapunov exponents it is not necessary to align with all of them to have a well-converged result, but Fig. 8(c) shows that the convergence times are, nevertheless, longer than the equipartition times shown in Fig. 2. Thus, the inescapable conclusion is that equipartition is reached before the trajectory has fully explored the chaotic sea.
It would be interesting to attempt a computation of the Lyapunov time of a submanifold, i.e. calculate $\tau_{1}$ starting from a nonthermal mode, e.g. $Q_{2}$, and see how this exponent relates to the true $\tau_{1}$ obtained from a thermal initial condition. Unfortunately this type of numerical investigation is hard, if not impossible, to carry on. This is due to the fact that the bushes associated to nonthermal modes in $\alpha$-FPUT chains are unstable and will not be preserved in long-time evolution no matter which numerical integrator is chosen [26]. Indeed, if the root chosen is nonthermal, by tracking the time-evolution of mode energies one sees that the initial excitation leaves the bush due to numerical errors and eventually covers the whole of mode space. This instability might be responsible for the observation that the evolution of nonthermal modes results in approximately the same Lyapunov times as thermal ones [23], which is surprising given that bush dynamics takes place only in a submanifold of mode space. Nevertheless, it is unclear how ergodicity in mode space relates to that of phase space, such that it is possible that a partial covering of mode space by a large enough bush is enough to resolve the “true” MLE. Besides, it might even happen that dynamics in phase space is ergodic while that in mode space is not, since ergodicity always depends on the observables chosen [41] unless it is tracked by means of invariant quantities [42]. At present, the authors are unaware of a special type of integrator that is capable of preserving bushes of excitations, such that studies regarding dynamical properties of chaotic subspaces cannot be performed in a numerically meaningful fashion. A consequence of this is that the energy in Fig. 1(a), which is correctly localized in the bush associated to $Q_{3}$, will eventually spread to all modes for long enough times. Nevertheless, the way “equipartition” takes place in this case is pathological and clearly traceable to numerical errors, as can be seen in Appendix A.
VI Conclusion
We have extended previous studies on the cubic Fermi-Pasta-Ulam-Tsingou ($\alpha$-FPUT) model by investigating the consequences of evolving higher-excited normal modes instead of the fundamental. The evolution of such atypical initial conditions was also compared to that obtained from typical (random) ones. We provided an explicit and simple way of predicting mode excitation sequences when exciting only a single normal mode (the root), showing that higher-excited roots take longer to cover mode space, while the fundamental is the fastest energy spreader. Thorough numerical investigations based on computing Lyapunov exponents and Kolmogorov-Sinai entropies show that even invariant quantities carry imprints of the atypical initial states used to compute them in their transient regime, despite converging to the correct invariant value for long times. The prethermal dynamics taking place before convergence is confirmed to be a consequence of some select roots lying close to stationary states also in the $\alpha$-FPUT system, and therefore being trapped in near-regular regions. The escape from such regions is what puts an end to the paradigmatic metastable plateaus in mode energies, whose end marks the moment invariant quantities computed from roots start to converge.
Acknowledgements.
This research was supported by the Institute for Basic Science through Project Code (No. IBS-R024-D1).
Appendix A Numerical instability of bushes
Figure 10: Examples of error-induced equipartition due to numerical instability of bushes. (a) Same as Fig. 1(a) but for longer times. (b) The evolution of $Q_{2}$, which as all even roots is nonthermal, in an $N{=}11$ chain. In both panels, states that are proven to be nonthermal end up reaching equipartition due to numerical errors. Panels (a) and (b) use different symplectic integrators (second and fourth orders), showing that the phenomenon happens independently of the solver, although it happens at different times depending on the time step (attesting for its numerical source).
The bushes associated with nonthermal roots in $\alpha$-FPUT chains appear to be unstable as they evolve in time. The main consequence of this instability is that energy, which should be only shared among modes in the bush and never leave for any finite time, will end up exciting other modes due to numerical errors. In Fig. 10 we provide two examples of this behavior when evolving provenly nonthermal modes: The roots $Q_{3}$ and $Q_{2}$ in a $N=11$ chain. The first case is essentially a continuation of Fig. 1 for much longer times, until one can see an anomalous and simultaneous jump in the energy of modes that should not be excited starting from $Q_{3}$; The root $Q_{2}$, on the other hand, is even and therefore always nonthermal, undergoing a very similar pathological jump and reaching equipartition in Fig. 10 exclusively due to numerical errors.